# Automated Static Vulnerability Detection via a Holistic Neuro-symbolic Approach

### v1

@@ Abstract.
Static vulnerability detection is still a challenging problem and demands excessive human efforts, e.g., manual curation of good vulnerability patterns. None of prior works, including classic program analysis or Large Language Model (LLM)-based approaches, have fully automated such vulnerability pattern generations with reasonable detection accuracy. In this paper, we design and implement, MoCQ, a novel holistic neuro-symbolic framework that combines the complementary strengths of LLMs and classical static analysis to enable scalable vulnerability detection. The key insight is that MoCQ leverages an LLM to automatically extract vulnerability patterns and translate them into detection queries, and then on static analysis to refine such queries in a feedback loop and eventually execute them for analyzing large codebases and mining vulnerabilities. We evaluate MoCQ on seven types of vulnerabilities spanning two programming languages. We found MoCQ-generated queries uncovered at least 12 patterns that were missed by experts. On a ground truth dataset, MoCQ achieved comparable precision and recall compared to expert-crafted queries. Moreover, MoCQ has identified seven previously unknown vulnerabilities in real-world applications, demonstrating its practical effectiveness. We have responsibly disclosed them to the corresponding developers.
@@ 1. Introduction
Static vulnerability detection examines source code without execution and thus is scalable and efficient for analyzing large codebases. It has been widely adopted in both industry and open-source security research. Classic static approaches conduct rigorous program analysis involving control- and data-flows on program representations (e.g., code property graph) generated from large codebases. One popular technique is to query the representations to search for given vulnerability patterns. This has been adopted by prominent tools like CodeQL (CodeQL, 2024), Semgrep (semgrep, 2025), and Joern (Yamaguchi et al., 2014; Backes et al., 2017), and has detected many zero-day vulnerabilities in the past (Backes et al., 2017; Yamaguchi et al., 2014; Shi et al., 2024; Shcherbakov et al., 2023).
Despite the success, one remaining barrier is that static vulnerability detection still requires excessive human efforts, especially manual curation of vulnerability patterns. Such manual efforts are not only time-consuming (e.g., spanning seven weeks (Fröberg, 2023)) but also introduce inaccuracies due to limited human knowledge, reducing the overall effectiveness of static detection (N/A, 2024; cod, 2024). As codebases evolve and new threats emerge, maintaining accurate detection patterns further requires continuous updates and efforts.
A natural solution to reduce the human efforts and inaccuracies is to introduce Large Language Models (LLMs). For example, researchers (Zhou et al., 2024; Lu et al., 2024; Ding et al., 2025) recently have explored using LLMs for static vulnerability detection without classic program analysis. The intuition is that LLMs excel at pattern recognition—they can identify recurring syntactic and semantic structures, and typical usage patterns, and understand semantics across diverse codebases and programming languages. For example, LLMs can understand programs based on natural language cues such as identifier names and comments. However, LLMs lack rigorous reasoning and often produce inconsistent or incomplete results due to their probabilistic nature and hallucinations. Additionally, token constraints limit their applicability to real-world complex software. Existing evaluation indicates that even state-of-the-art LLMs performed poorly with high false positive rates and often non-deterministic outputs (Ullah et al., 2024).
@@
Since neither the classic nor LLM-based static approach is sufficient for effective vulnerability detection, a promising direction is to integrate them into a neuro-symbolic framework [^1] Neuro refers to LLM-based analysis, which leverages neural models for semantic reasoning; and symbolic refers to classic static analysis with symbolic representations and formal logic systems. that combines their complementary strengths. Some recent works have already taken a step forward in such an integration. For example, IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) use LLMs to extract taint specifications such as sources and sinks and integrate them into existing detection patterns. However, human efforts are still needed, indicating significant engineering efforts and inaccuracies due to human errors also exist because they rely on existing expert-crafted detection patterns.
@@
In this paper, we design a novel, static, fully automated, neuro-symbolic vulnerability detection framework, called MoCQ (Model-Generated Code Queries), which splits the tasks for LLMs and static program analysis based on their respective strengths. The core idea is to leverage LLMs to automatically extract vulnerability patterns and incorporate the patterns into static program analysis. This achieves the best of both worlds: the pattern recognition capability of LLMs, and the scalability and rigorousness of classic static analysis. More specifically, MoCQ uses LLMs to generate vulnerability patterns and express them as queries to the program representations generated via classic program analysis.
@@
While the idea is intuitively simple, the LLM-driven generation of vulnerability patterns faces multiple challenges. LLMs often do not have enough knowledge of domain-specific languages (DSLs) used by static vulnerability detection tools (e.g., Scala-like DSL for Joern (Yamaguchi et al., 2014) and SQL-like DSL for CodeQL), resulting in unparseable or uncompilable queries. Specifically, our study shows that pretrained, commercial LLMs, e.g., ChatGPT (OpenAI, 2024), often fail to generate valid queries, frequently producing syntax errors or triggering execution-time exceptions. To deal with this challenge, MoCQ adopts a technique, specifically DSL subsetting, to refine DSL by selecting a core set of features to guide the automated query generation, without sacrificing expressiveness. The intuition is that such a refined subset can be better understood by pretrained LLMs while still preserving the core semantics due to the redundant nature of the DSL.
@@
Besides, even when the LLM-generated queries are syntactically correct and executable, they may fail to capture vulnerabilities effectively. Specifically, they may trigger semantic errors (i.e., fail to report vulnerabilities), be overly specific (i.e., overfitting to known instances via matching exact variable names or fixed control flow paths), or be overly general (i.e., underfitting by matching many irrelevant code cases). To deal with these, MoCQ adopts a feedback-driven approach, which incorporates fine-grained feedback from a symbolic query validator for iterative query refinement. The query validator executes the LLM-generated queries and monitors the exhibited behaviors, including syntax errors, execution exceptions, and semantic errors. Specifically, we instrument the query execution runtime to obtain a block-level program state of the query execution to precisely locate the errors and inconsistencies. These are then used for the LLM to refine the incorrect query. The validator further uses heuristics to detect if a query is too specific and instructs the LLM to generalize it; it also instructs LLMs to eliminate false positives (if any) to improve the precision.
@@
We implemented MoCQ for two state-of-the-art static vulnerability detection tools (Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024)). We then extensively evaluated MoCQ on seven types of vulnerabilities from the OWASP Top Ten (OWASP Foundation, 2021), covering two popular programming languages PHP and JavaScript. Our results show that, given only a few (e.g., 5-18) vulnerability examples, MoCQ can efficiently generate detection queries within hours. These queries demonstrate strong detection capability, comparable to those crafted by security experts. Notably, MoCQ uncovered at least 12 vulnerability patterns that were missed in expert-crafted queries, highlighting the strength of our LLM-based approach. A comprehensive ablation study further confirms the importance of DSL subsetting and the symbolic validator, which significantly contribute to the validity of the generated queries. MoCQ-generated queries successfully detected seven new vulnerabilities in real-world applications and four could not be identified by expert-crafted patterns.
@@
  • We developed a DSL subsetting technique and a feedback-driven approach to automatically generating vulnerability patterns.
  • We implemented a holistic, fully-automated, neuro-symbolic vulnerability detection framework MoCQ.
@@
  • We demonstrated that MoCQ could automatically generate vulnerability patterns for real-world vulnerabilities and achieve comparable results with patterns from security experts. MoCQ uncovered at least 12 new patterns and seven new vulnerabilities.
2. Background
2.1. Large Language Models
Recent advancements in LLMs such as ChatGPT (OpenAI, 2024), Claude (Anthropic, 2025), and Gemini (DeepMind, 2023) have demonstrated their capability in various tasks including code generation, unit testing, and reasoning over structured data (Du et al., 2024; Touvron et al., 2023). Trained on massive codebases, these models can understand complex programming structures by learning the cues from natural language semantics such as identifier names and human-written comments. For example, Artemis (Ji et al., 2025) uses an LLM to analyze the function name to decide if a function accepts user inputs in web applications. LLMs can also generate syntactically and semantically correct code snippets to solve code challenges like LeetCode. Some attempts leverage LLMs for security-relevant tasks. For example, Fuzz4All (Xia et al., 2024) uses LLMs to generate fuzzing inputs; ChatAFL (Meng et al., 2024) employs LLMs to extract input format (grammar) from network protocol specifications for protocol fuzzing. These efforts showcase the effectiveness of LLMs in extracting information from natural language descriptions and code.
[…]
In-context Learning.In-context LearningIn-context Learning. In-context learning is a fundamental capability of modern LLMs, allowing them to perform new tasks without explicit retraining or post-training. Instead of modifying model weights, in-context learning enables LLMs to generalize patterns and behaviors from a few examples as part of input prompts like in few-shot learning.
2.2. Query-based Static Analysis
Query-based static analysis tools formulate vulnerability patterns as structured queries in the domain-specific languages (DSLs) of the tools. They typically employ a language-specific frontend to convert source code into graph-like program representations, such as an abstract syntax tree (AST), control-flow graph (CFG), or code property graph (CPG) (Yamaguchi et al., 2014). Security analysts then define precise analysis queries that operate on the representations to find vulnerabilities. By decoupling analysis logic (queries) from the language semantic reasoning backend, this approach could achieve high flexibility and maintainability. For example, to extend to a different programming language, one only needs to develop an additional language frontend and could reuse the backend. Notable tools include Joern (Yamaguchi et al., 2014; Backes et al., 2017), GitHub’s CodeQL (CodeQL, 2024), and Amazon’s CodeGuru (Services, 2025). More specifically, Yamaguchi et al. introduced Joern by modeling programs in CPG (joe, 2025). GitHub’s CodeQL (CodeQL, 2024) further popularized query-based analysis by integrating it into security workflows for large-scale vulnerability detection.
The query DSLs allow security analysts to specify vulnerability patterns. Each analysis tool typically comes with its own DSL. For example, Joern uses a Scala-like DSL with a rich set of APIs and custom data structures. CodeQL, on the other hand, adopts a SQL-like DSL. During query execution, the static analysis tools often enforce strict syntax and semantic validations.
3. Problem Statement
3.1. Motivation
The effectiveness of query-based static analysis hinges not only on rigorous language semantic reasoning but more importantly, on the quality of vulnerability patterns. High-quality, well-crafted patterns are essential. Without them, even the most powerful static analysis may miss critical vulnerabilities or generate excessive noise.
@@ 3.1. Motivation
1// find obj["__proto__"] in t = obj["__proto__"] 2def objProto = cpg.call3.where(_.name(Operators.assignment))4.argument(2)5.isCall6.arrayAccess78// find t["toString"] in t["toString"] = "Hacked"9def objProp = cpg.call10.where(_.name(Operators.assignment))11.argument(1)12.isCall13.arrayAccess1415// find a data-flow path from objProp to objProto16objProp.array.reachableBy(17    objProto18)19.filter(...)
@@
[…]
A Query Example.A Query ExampleA Query Example. Figure 1 illustrates a simplified version of a real-world Joern query designed to JavaScript detect prototype pollution (Fröberg, 2023). Specifically, JavaScript is a prototype-based language where objects inherit properties and methods through a mechanism known as prototypes. Modifying the prototype of one object can affect all other objects that inherit from the same prototype, resulting in vulnerabilities known as prototype pollution. For example, an attacker can overwrite the toString method by first accessing the prototype reference using t=obj["__proto__"] and then modifying it via t["toString"]="Hacked". As a result, any object inheriting from the fundamental prototype Object.prototype now have a tampered toString method. This type of vulnerability can lead to severe security consequences, including arbitrary code execution, cross-site scripting, privilege escalation, and denial of service (Shcherbakov et al., 2023; Li et al., 2021a; Cornelissen et al., 2024; Liu et al., 2024a; Kang et al., 2022a).
Figure 1 comprises three main steps: (1) identifying the object prototype (lines 1-6), (2) identifying the property access (lines 8-13), and (3) verifying whether there is a data connection between them (lines 15-19). Constructing such queries requires a profound understanding of the query DSL and is inherently challenging. [^2] The original query is more complex and includes additional operations, such as verifying whether the accessed values and properties are controllable by user inputs. We simplified it for illustration and clarification purposes. As an example, consider the first step. The dot operator (".") in Joern’s DSL enables query chaining, meaning that each stage operates on the intermediate results produced by the preceding stage to progressively refine the search. The query begins by identifying assignment operations via Operators.assignment (in Joern, assignment is modeled as a form of call operation). It then extracts the second argument of the assignment using argument(2), which corresponds to the right-hand side (RHS) of the assignment expression. Finally, arrayAccess is used to filter results, narrowing the matches to cases where the RHS is an indexed (array-style) access.
[…]
Summary.SummarySummary. Designing robust vulnerability patterns remains a significant challenge. First, it requires expertise in security vulnerabilities, programming languages, the static analysis tool, and DSLs, resulting in a time-consuming process that is difficult to get right. Maintaining these queries to account for emerging threats further demands ongoing expertise and sustained efforts. For example, from the Git history, the version shown in Figure 1 took three weeks to develop, followed by an additional month of refinement (que, 2023). Second, experts cannot always capture the full breadth of attack vectors, leading to missed vulnerabilities. In practice, many queries—even those written by GitHub’s CodeQL team—have been found to be imprecise or buggy (N/A, 2024; cod, 2024). In the above example, an attacker can bypass the query detection by chaining property access and assignment in a single expression, such as obj["__proto__"]["toString"] = "Hacked". Beyond that, as we will show in § 6.3.2, three more advanced mechanisms for manipulating object properties—automatically uncovered by our research—were overlooked even in the latest version of the example query.
@@ 3.2. Challenges
This work aims to explore an automated approach to generating vulnerability patterns using LLMs. When new threats are publicly disclosed, a developer can rapidly generate comprehensive vulnerability patterns to scan their codebases without the long delays traditionally required for manual pattern engineering. While the idea of using LLMs to generate queries is intuitively simple, the task is challenging as shown in our study on five vulnerability types among the OWASP Top Ten (OWASP Foundation, 2021). Specifically, we asked ChatGPT-4o (OpenAI, 2024) to generate vulnerability queries for these vulnerabilities. Among all 300 cases spanning five vulnerability types with three prompting strategies, only one successfully generated a query to retrieve a JavaScript command injection under the few-shot setting. More details can be found in Appendix A.
(Simplified) Prompt : You are an expert on Joern and static analysis. You are given two query examples and a vulnerability example, please generate a Joern query to detect JavaScript prototype pollution. (Simplified) Answer: 1... 2val assign = code.call 3  .nameExact("assignment") 4  ...
In the above, we show a prompt to generate a query for JavaScript prototype pollution under the few-shot setting, and the response query. The generated query actually could not successfully run in Joern’s analysis engine. We summarize the following four key technical problems.
[…]
@@
DSL Syntax Correctness.DSL Syntax CorrectnessDSL Syntax Correctness. The LLM-generated queries frequently raise grammar and syntax errors, which stem from the model’s incomplete understanding of the DSL grammar. Specifically, static analysis tools use complex DSLs, and LLMs typically lack sufficient exposure to their syntax, semantics, and typical usage. For example, Joern queries based on the code property graph usually require the root object cpg (e.g., Figure 1). However, the query generated above did not follow this requirement. This issue is further aggravated by the hallucinations of LLMs.
[…]
@@
Runtime Execution Correctness.Runtime Execution CorrectnessRuntime Execution Correctness. Queries must execute within domain-specific runtime (i.e., analysis engine) and interact meaningfully with structured program representations. LLM-generated queries often use incorrect or non-existent function calls, missing parameters, and incompatible data types or operations, resulting in execution-time exceptions. For example, the nameExact operation in the above query used "assignment" instead of "<operator>.assignment", which led to execution-time exceptions after we first manually fixed the syntax error.
[…]
@@
Semantic Validity.Semantic ValiditySemantic Validity. Even when queries are syntactically correct and executable, ensuring their semantic validity remains challenging. They must satisfy runtime constraints and capture the correct traversal logic. Executable queries may still fail to reflect intended program behaviors, leading to semantic errors, where the query runs without crashing but misses the vulnerable code.
[…]
@@
Precision-recall Balance.Precision-recall BalancePrecision-recall Balance. Similar to how security analysts craft queries, LLM-generated queries must be not only correct but also comprehensive and precise. Patterns that are too specific (i.e., overfitting) may miss variants of the same vulnerability, resulting in high false negatives, whereas overly general patterns (i.e., underfitting) can produce excessive false positives. Achieving this balance between precision and recall remains a fundamental challenge in query generation.

### v3

@@ Abstract.
In this paper, we present MoCQ, a novel neuro-symbolic framework that combines the complementary strengths of Large Language Model (LLM) and classic vulnerability checkers to enable scalable, automated vulnerability detection. The key insight is to leverage an LLM to automatically generate vulnerability patterns and translate them into detection queries. Specifically, MoCQ incorporates an iterative loop in which an LLM refines queries based on carefully designed feedback information. The resulting queries are then executed to analyze large codebases and detect vulnerabilities. We evaluated MoCQ on 12 vulnerability types across four programming languages. MoCQ achieved comparable precision and recall compared to expert-developed queries, with significantly less expert time needed. MoCQ also uncovered 46 new vulnerability patterns that experts missed, each representing an overlooked vulnerability class. MoCQ further discovered seven previously unknown vulnerabilities in real-world applications.
@@ 1. Introduction
Pattern-based static analysis (i.e., vulnerability checker) identifies vulnerabilities by matching source code against known vulnerable patterns, without requiring program execution. Compared to dynamic techniques (Pham et al., 2017; Trickel et al., 2023; Lee et al., 2021), it scalably analyzes large codebases with high code coverage, and requires less setup effort (e.g., no need for software deployment). It has been widely adopted in practice (Backes et al., 2017; Yamaguchi et al., 2014; Shi et al., 2024; Shcherbakov et al., 2023) and has discovered numerous high-impact vulnerabilities (e.g., Log4Shell (Services, 2021; MITRE Corporation, 2021)). For instance, CodeQL (CodeQL, 2024a) is extensively integrated into GitHub’s CI/CD pipelines and is employed to analyze large-scale applications such as the Chrome browser (Project, 2025). Joern (Yamaguchi et al., 2014; Backes et al., 2017) has been recognized as one of the most widely used tools for automated vulnerability discovery (Shiri Harzevili et al., 2024).
Despite its success, the effectiveness of this approach hinges entirely on the quality of vulnerability patterns. These patterns describe vulnerable code structures or behaviors and are used to match new vulnerability instances in target codebases. However, the creation of these patterns is largely manual and demands substantial human effort. This process is not only time-consuming (e.g., several weeks or more (Fröberg, 2023)), but also prone to inaccuracies due to limited human knowledge. This significantly reduces the effectiveness of detection (N/A, 2024; cod, 2024). As codebases evolve and new threats emerge, keeping detection patterns accurate and up to date requires ongoing maintenance and engineering effort.
To alleviate this burden, researchers have alternatively explored learning-based approaches that analyze source code without execution. Most deep learning (DL)-based methods operate at the function level to classify functions as vulnerable or not. However, they often lack interprocedural or project-wide context (Li et al., 2021b; Russell et al., 2018), and require a large volume of labeled vulnerabilities for training (Chakraborty et al., 2021), which may not be available. More recently, several works (Zhou et al., 2024; Lu et al., 2024; Ding et al., 2025) have used Large Language Models (LLMs) to directly analyze code without relying on traditional program analysis. Leveraging massive training corpora, LLMs can recognize typical (vulnerable) usage patterns and reason about code semantics across diverse programming contexts. While promising, LLMs face challenges such as token length limits, inconsistent reasoning, and susceptibility to hallucinations, which limit their reliability in analyzing large, complex software.
@@
The distinct challenges of classic and learning-based methods suggest that a hybrid neuro-symbolic approach [^1] Neuro refers to learning-based analysis (particularly LLMs in this paper) leveraging neural models for semantic reasoning; symbolic refers to classic static analysis using symbolic representations and formal logic systems. is a promising direction. While recent works have taken steps toward such integration, they fall short of the level of automation. For example, IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) use LLMs to extract taint sources and sinks, and incorporate them into existing detection patterns to enhance the coverage of vulnerability searches. However, these approaches only extend some parts of base patterns developed by humans, which still require manual effort and are prone to human error.
@@
In this paper, we design the first static, automated neuro-symbolic vulnerability detection framework, called MoCQ (Model-Generated Code Queries), that combines the complementary strengths of LLMs and classic static analysis. The core idea is to leverage LLMs to automatically generate end-to-end vulnerability patterns and incorporate the patterns into static program analysis. Such a neuro-symbolic design not only preserves the scalability and rigorousness of classic static analysis, but also gets rid of the time-consuming pattern engineering.
@@
However, LLM-driven generation of vulnerability patterns faces multiple challenges. LLMs often do not have enough knowledge of domain-specific languages (DSLs) used by static vulnerability detection tools (e.g., Scala-like DSL for Joern (Yamaguchi et al., 2014) and SQL-like DSL for CodeQL), resulting in uncompilable queries. Specifically, our study shows that pretrained, commercial LLMs, e.g., GPT-4o (OpenAI, 2024), often fail to generate valid queries, frequently producing syntax errors or triggering execution-time exceptions. To deal with this challenge, MoCQ adopts a technique, specifically DSL subsetting, to refine DSL by selecting a core set of features to guide the automated query generation, without sacrificing expressiveness. The intuition is that such a refined subset can be better understood by pretrained LLMs while still preserving the core semantics due to the redundant nature of the DSL.
@@
Even when the LLM-generated queries are syntactically correct and executable, they may fail to capture vulnerabilities effectively. Specifically, they may trigger semantic errors (i.e., fail to report vulnerabilities), be overly specific (i.e., overfitting to known instances via matching exact variable names or fixed control flow paths), or be overly general (i.e., underfitting by matching many irrelevant code cases). To deal with these, MoCQ adopts a feedback-driven approach, which incorporates fine-grained feedback from a symbolic query validator for iterative query refinement. The query validator executes the LLM-generated queries and monitors the exhibited behaviors, including syntax errors, execution exceptions, and semantic errors. Specifically, we instrument the query execution runtime to obtain a block-level program state of the query execution to precisely locate the errors and inconsistencies. These are then used for the LLM to refine the incorrect query. The validator further uses heuristics to detect if a query is too specific and instructs the LLM to generalize it; it also instructs LLMs to eliminate false positives (if any) to improve the precision.
@@
We implemented MoCQ for two state-of-the-art static detection tools, Joern (Yamaguchi et al., 2014) and CodeQL (CodeQL, 2024a). MoCQ is language-agnostic and can generate patterns for all programming languages supported by these tools. We extensively evaluated MoCQ on 12 representative vulnerability types (e.g., use-after-free and cross-site scripting) on four popular programming languages. Our results show that, given only a few (e.g., 5-18) vulnerability examples, MoCQ can efficiently generate detection queries within hours. These queries demonstrate strong detection capability comparable to that developed by security experts. MoCQ-generated queries successfully detected seven new vulnerabilities in real-world applications.
Beyond detection performance, MoCQ uncovered 46 new vulnerability patterns missed by expert-developed queries, with five patterns already integrated into upstream tool repositories. Each missing pattern often results in overlooking entire vulnerability classes. Our further analysis revealed complementary strengths of LLMs and human experts: LLMs excel at capturing subtle linguistic nuances that experts consistently overlook, while experts demonstrate superior high-level reasoning for complex vulnerabilities.
@@
  • We developed a DSL subsetting technique and a feedback-driven approach in MoCQ, which automatically generates end-to-end vulnerability patterns.
@@
  • We demonstrated that MoCQ could automatically generate vulnerability patterns for real-world vulnerabilities and achieve comparable results with patterns from security experts.
  • MoCQ uncovered 46 new patterns and seven new vulnerabilities.
2. Background and Motivation
2.1. Pattern-based Static Analysis
Pattern-based static analysis is a representative and practical approach among various static analysis techniques. Tools (e.g., Joern (Yamaguchi et al., 2014; Backes et al., 2017) and CodeQL (CodeQL, 2024a)) express vulnerability patterns as structured queries written in tools’ domain-specific languages (DSLs). They use language-specific frontends to translate source code into graph-like program representations such as control-flow graphs (CFG) and code property graphs (CPG) (Yamaguchi et al., 2014). Security analysts then write precise queries that operate on these representations to identify vulnerabilities. Joern models programs with CPGs (joe, 2025), while CodeQL relies on code databases. The DSLs differ as well: Joern employs a Scala-like language, whereas CodeQL uses SQL-like syntax. Static analyzers enforce strict syntax and semantic checks during query execution to ensure accuracy and soundness.
Example: A Prototype Pollution Query. JavaScript is a prototype-based language where objects inherit properties and methods through a mechanism known as prototypes. Modifying the prototype of one object can affect all other objects that inherit from the same prototype, resulting in prototype pollution. For example, an attacker can access the prototype via p=obj["__proto__"] and modify with it using p["toString"]="Hacked". As a result, any object inheriting from the fundamental prototype Object.prototype now has tampered toString method. This type of vulnerability can lead to arbitrary code execution, privilege escalation, and denial of service (Shcherbakov et al., 2023; Li et al., 2021a; Cornelissen et al., 2024; Liu et al., 2024; Kang et al., 2022).
@@ 2.1. Pattern-based Static Analysis
  1 // find obj["__proto__"] in p=obj["__proto__"]  2 def objProto = cpg.call  3   .where(_.name(Operators.assignment)).argument(2).isCall.arrayAccess   4   5 // find p["toString"] in p["toString"]="Hacked"  6 def objProp = cpg.call  7   .where(_.name(Operators.assignment)).argument(1).isCall.arrayAccess   8   9 // find a data-flow path from objProp to objProto  10 objProp.array.reachableBy(objProto).filter(...)
@@
Figure 1 shows a simplified version of a real-world Joern query designed to detect JavaScript prototype pollution (Fröberg, 2023). [^2] The original query is more complex, including checks for user-controlled values, property filtering, etc. It comprises three main steps: (1) identifying the object prototype (lines 1-3), (2) identifying the property access (lines 5-7), and (3) verifying whether there is a data connection between them (lines 9-10). For example, step (1) uses Joern’s dot operator (".") to chain operations, where each stage filters results produced by the previous one. It begins by selecting assignment expressions via Operators.assignment, [^3] In Joern, assignment is modeled as a form of call operation. extracts the right-hand side via argument(2), and narrows results to array-style accesses using arrayAccess().
Query Analysis. Writing high-quality patterns is non-trivial. For example, the query in Figure 1 took about three weeks to draft and an additional month to refine (que, 2023). Experts may miss attack variants or write overly general patterns, causing false negatives and false positives (N/A, 2024; cod, 2024). For instance, attackers can bypass the detection by combining property access and assignment in one expression, like obj["proto"]["toString"]="Hacked". This also happens in the official CodeQL queries maintained by GitHub Security.
This work thus explores an automated approach to generating vulnerability patterns using LLMs. When a new threat is publicly disclosed, a developer can rapidly generate comprehensive vulnerability patterns to scan their codebases, without the long delays traditionally required for manual pattern engineering.
@@ 2.2. Challenges
While using LLMs to generate queries may seem intuitive, our preliminary study uncovers substantial challenges. We tasked GPT-4o (OpenAI, 2024) with generating vulnerability queries for five types of vulnerabilities using three prompting strategies: zero-shot, few-shot, and chain-of-thought (CoT). In each case, we provided an input vulnerability example as the detection goal. Out of 300 cases, only one successful query was produced in the few-shot setting to retrieve the input vulnerability example. We omit full results due to space constraints. We summarize four key technical challenges. Below, we show a LLM-generated query for JavaScript prototype pollution that failed to execute in Joern’s analysis engine.
An invalid query generated in few-shot prompting:   1 ...   2 val assign = code.call // omitting ‘cpg‘   3   .nameExact("assignment") // using wrong operator name
@@
DSL Syntax Correctness. The LLM-generated queries frequently raise grammar and syntax errors, which stem from the model’s incomplete understanding of the DSL grammar. Specifically, LLMs typically lack sufficient exposure to the syntax, semantics, and typical usage of complex DSLs. For example, Joern queries based on the code property graph usually require the root object cpg. However, the query generated above did not follow this requirement. This issue is further aggravated by the hallucinations of LLMs.
@@
Runtime Execution Correctness. Queries must execute within the domain-specific runtime (i.e., analysis engine) and interact meaningfully with structured program representations. LLM-generated queries often use incorrect or non-existent function calls, missing parameters, and incompatible data types or operations, resulting in execution-time exceptions. For example, the nameExact operation in the above query used "assignment" instead of "<operator>.assignment", which led to execution-time exceptions after we first manually fixed the syntax error.
@@
Semantic Validity. Even when queries are syntactically correct and executable, ensuring their semantic validity remains challenging. They must satisfy runtime constraints and capture the correct traversal logic. Executable queries may still fail to reflect intended program behaviors, leading to semantic errors, where the query runs without crashing but misses the vulnerable code.
@@
Precision-recall Balance. Similar to how security analysts develop queries, LLM-generated queries must be not only correct but also comprehensive and precise. Patterns that are too specific (i.e., overfitting) may miss variants of the same vulnerability, resulting in high false negatives, whereas overly general patterns (i.e., underfitting) can produce excessive false positives. Achieving this balance between precision and recall remains a fundamental challenge.